OpenStudy (anonymous):

Find cot θ if csc θ = square root 5 / 2 and tan θ > 0

OpenStudy (anonymous):

\[\csc \theta~is~positive.~so~\theta ~lies~\in~1st~and~2rd~quadrant.\] \[\tan \theta>0,~so~\theta~lies~\in~1st~and~3rd~quadrant.\ Hence cot theta is positive. \[\csc ^2\theta-\cot ^2\theta=1\] find cot theta

OpenStudy (triciaal):